OpenStudy (anonymous):

An airplane flew 3 hours with a 60 mph head wind. The return trip with a tail wind of the same speed took 2 hours. Find the speed of the plane in still air. A. 380 mph B. 340 mph C. 320 mph D. 300 mph

OpenStudy (anonymous):

head wind, REDUCES the planes speed, so: rate 1 = (x-60) mph ; time = 3hrs rate 2 = (x+60) mph ; time = 2hrs the distances are the same and distance is rate *time ,hence: (x-60) * 3 = (x+60) * 2 can you proceed from here?

OpenStudy (anonymous):

x=300

OpenStudy (anonymous):

right